Problem 10.38 Part A A neon sign is made of glass tubing whose inside diameter is 2.5 cm and whose length is 5.5 m If the sign contains neon at a pressure of 1.77 torn at 31 °C, how many grams of neon are in the sign? (The volume of a cylinder is rh) Express your answer using two significant figures. In Asia on a , 7 Submit My Answers Give Up Continue Part The laser pointer emits light because electrons in the material are excited (by a battery) from their ground state to an upper excited state. When the electrons return to the ground state, they lose the excess energy in the form of 630 nm photons. What is the energy gap between the ground state and excited state in the laser materia? VO AEs n o B, 9 AE Submit My Answers Give Up Atoms involved in bonds are usually more stable than they would be otherwise. This means that energy is released in the formation of the bond For ionic bonds, Coulombs law is used to calculate this energy Submit Hints My Answers Give Up Review Pant Correct Part B where E is energy in joules J. k is a constant equal to 2.31 x 10 J. m , and an are the charges of the two ions, and d is the distance between the nuclei of the two lons in meters (m). Arrange the following ionic compounds in order of decreasing amount of energy released in lattice formation: NaFBao, InN, and KBr Whereas the bond energy calculated by Coulombs law pertains to a single ionic bond, lattice energy pertains to all the ionic bonding among a group of atoms arranged in a crystal lattice. Lattice energy is calculated by using the Coulombs law equation, but with a different constant (unique to each substance) that takes into account the crystalline structure of the substance. Rank from greatest to least lattice energy. To rank items as equivalent, overlap them. Naf Bao InNKBr By convention, lattice energy is defined as the amount required to either break an onic solid into individual gaseous ions, or to form an ionic solid from gaseous ions. Greatest SmallestPlease help out on any of these! Thank you!!
Answer #1

Part A

radius of tube = 2.5 / 2 cm = 0.0125 m

length of tube = 5.5 m

Volume of tube = \pir2h

= 2.699x10-3 m3

= 2.699 L

Pressure = 1.77 torr

= 0.0023 atm

temperature = 31 oC

= 304 K

Using PV = nRT

0.0023*2.699 = n*0.082*304

0.00620 = n*24.928

n = 0.000248 moles

molar mass of neon = 20.17 g/mol

mass of neon gas = 20.17*0.000248

= 5.0x10-3 g

Part B -

Decreasing order - NaF > KBr > InN > BaO